peter can crack the can in 30 minutes. renato can crack he can in 40 minutes. after working together for 20 minutes, badong helped them and finish the job in 2 minutes. if badong will do the entire job alone, for how long will it take him to finish the job?
Answer by richwmiller(17219)

The problem is peter and renato already finished the job in less than 18 minutes.
a=30
b=40
1/a+1/b=1/x
1/30+1/40=1/x
1/30+1/40=1/17.1428571
1/x=1/17.1428571
x=17.1428571 minutes working together
check
1/a+1/b=1/x
1/30+1/40=1/x
40/1200+30/1200=1/x
70/1200=1/x
1200/70=x
x=17.1428571
ok